OpenStudy (anonymous):

What is the x-intercept and y-intercept of the graph of 10x - 5y = 20?

OpenStudy (anonymous):

x-intercept = (0, -4); y-intercept = (2, 0) x-intercept = (0, 4); y-intercept = (-2, 0) x-intercept = (-2, 0); y-intercept = (0, 4) x-intercept = (2, 0); y-intercept = (0, -4)

OpenStudy (anonymous):

What grade are you in?

OpenStudy (anonymous):

Can I give you a hint? What is y going to be when the line intercepts x?

OpenStudy (anonymous):

*intersects

OpenStudy (anonymous):

x-intercept = (0, -4); y-intercept = (2, 0)

OpenStudy (anonymous):

II mean x-intercept = (2, 0); y-intercept = (0, -4)

OpenStudy (anonymous):

Nullify needs to try some of these for himself, lol. That's the only way you learn.